### v2

# BARE: Combining Base and Instruction-Tuned Language Models for Better Synthetic Data Generation

Abstract
As the demand for high-quality data in model training grows, researchers and developers are increasingly generating synthetic data to tune and train LLMs. A common assumption about synthetic data is that sampling from instruct-tuned models is sufficient; however, these models struggle to produce diverse outputs—a key requirement for generalization. Despite various prompting methods, in this work we show that achieving meaningful diversity from instruct-tuned models remains challenging. In contrast, we find base models without post-training exhibit greater diversity, but are less capable at instruction following and hence of lower quality. Leveraging this insight, we propose Base-Refine (BARE), a synthetic data generation method that combines the diversity of base models with the quality of instruct-tuned models through a two-stage process. With minimal few-shot examples and curation, BARE generates diverse and high-quality datasets, improving downstream task performance. We show that fine-tuning with as few as 1,000 BARE-generated samples can reach performance comparable to the best similarly sized models on LiveCodeBench tasks. Furthermore, fine-tuning with BARE-generated data achieves a 101% improvement over instruct-only data on GSM8K and a 18.4% improvement over SOTA methods on RAFT.
Machine Learning, ICML
Figure 1: Histogram of pairwise embedding similarities for 1000 Llama-3.1-70B-Base vs Instruct generations of grade school math problems. The base distribution is further to the left, indicating lower similarity and hence higher diversity. Note that frequency is of the cosine similarity score.
Figure 2: Accuracy of a Llama-3.2-1B-Instruct model fine-tuned on 4 different sets of synthetically generated math problems, evaluated on a randomly selected $n=500$ subset of GSM8K. Training with BARE-generated data outperforms both Base and Instruct generations.
1 Introduction
As Large Language Models (LLMs) grow in size and capability, the demand for high-quality, diverse data in model training is outpacing human-generated data, necessitating the use of synthetically generated data (Villalobos et al., 2024). Consequently, we expect the compute spent on curating and generating data for model training to increase significantly in the coming years, especially in low-data domains where high-quality data is scarce. For example, it is common to use LLMs to generate synthetic data for a variety of tasks such as math, code, function calling, general reasoning, etc. (Yu et al., 2024; Guo et al., 2024; Patil et al., 2023; Liu et al., 2024).
The ease of data generation has led many dataset creators (Samvelyan et al., 2024; Zhang et al., 2024c; NovaSky Team, 2025) and model creators (Dubey et al., 2024; Qwen et al., 2025; Nvidia et al., 2024; Guan et al., 2025; NovaSky Team, 2025) to turn to synthetic data sampled from instruct-tuned language models as a replacement for tasks where human-generated data is lacking.
Challenges with Instruct-Tuned Models.
Though instruct-tuned models have become more and more capable at complex tasks resulting in higher quality generations, there have been discussions of how accepted methods for post-training can lead to mode collapse (Shumailov et al., 2024; Wong et al., 2024; Lambert et al., 2024). Mode collapse refers to an inability for models to generate diverse outputs to queries that don’t have a single response, yet synthetic data is primarily useful when it is both high quality and diverse (Chen et al., 2024; Raventós et al., 2023).
Intuitively, one might expect a variety of prompting techniques that explicitly encourage diversity to address this gap. For example, prompting methods we baseline against include conditioning a generation on all previous ones and prompting for it to be distinct, assuming different personas for different generations, or asking for multiple distinct entries in a single call (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024). However, we find that the amount of meaningful diversity from such methods is limited and cannot fully address the gap introduced by post-training. On an open-ended generation task like email generation, such prompting methods do not meaningfully improve dataset diversity (Table 2) or trained model accuracy (Figure 9).
Base vs. Instruct-Tuned Models.
An alternative approach is to leverage base models, which are not constrained by post-training biases and thus produce outputs that better represent diversity present in real-world data (OpenAI, 2024b). More quantitatively, as shown in Figure 2, base models generate outputs with noticeably lower pairwise cosine similarity (mean similarity: $0.313$) compared to instruct-tuned models (mean similarity: $0.421$). Here we use cosine similarity as a quantitative surrogate for diversity — a dataset with greater cosine similarity has more similar items and therefore less diversity. The improved diversity in base model generations can improve downstream task performance, with Figure 2 showing that the accuracy of Llama-3.2-1B-Instruct on GSM8K increases to 22.5% from 17.8% when fine-tuning on the Llama-3.1-70B-Base generated dataset of word problems rather than the Llama-3.1-70B-Instruct generated dataset.
BARE.
In this work, we observe that diversity is critical for synthetic data generation when used in downstream training and fine-tuning tasks. We find that while the data generated from a base model may be more diverse, it also tends to be of lower quality (Figure 5). This can counteract the gains from diversity and hinder downstream training. Our key insight is that by combining base and instruct-tuned models, one can improve the diversity of a dataset while controlling the quality of individual data entries.
To this end, we introduce Base-Refine (BARE) — a novel approach that leverages the diverse data generated by the base models and refines it with instruct-tuned models. In a variety of settings, we find this two-stage process enhances diversity without compromising quality, enabling the generation of datasets that improve downstream performance.
We evaluate fine-tuning with BARE-generated data on contemporary tasks, such as the recently introduced Retrieval-Augmented Fine-Tuning (RAFT) method (Zhang et al., 2024c). RAFT generates synthetic question/answer pairs starting from just a corpus to fine-tune a RAG model, but we find that the diversity of data it generates is low. By replacing the data generator in RAFT with BARE, we show we can achieve up to a 18.4% fine-tuned accuracy lift over the existing work that solely uses instruct-tuned models.
As another example, BARE improves the generation of synthetic math training data. We generated grade school math problems similar to GSM8K problems with several methods, shown in Figure 2. We find that refining Llama-3.1-70B-Base data generations with Llama-3.1-70B-Instruct improves a fine-tuned model’s accuracy on GSM8K from 22.4% (base) to 29.8% (refined). Likewise, using GPT-4o as a generator out of the box (22.4%) does not perform as well as using GPT-4o as a refiner of generations from Llama-3.1-70B-Base (35.8%).
To summarize, our contributions are:
  1. 1.
    We quantitatively investigate the quality and diversity of base and instruct-tuned models for synthetic data generation across various sampling methods to motivate better system design. In Section 3, we find that base models tend to produce more diverse responses whereas instruct-tuned models offer higher quality.
  2. 2.
    Using these insights, in Section 4, we propose Base-Refine (BARE), a practical new method for generating synthetic data. We show in Section 5 that BARE consistently improves fine-tuned model performance over various baselines, including SOTA generation methods, and with only 1000 samples can lead to fine-tuned models comparable to SOTA models of similar sizes.
2 Related Work
Synthetic Data.
Synthetic data is commonly used to train state of the art models like Llama 3.3 (Dubey et al., 2024), Qwen 2.5 (Qwen et al., 2025), Nemotron 4 (Nvidia et al., 2024), and o3-mini (Guan et al., 2025). However, prior work has shown its usage poses risks such as model collapse, where iterative training on low-diversity data shifts the generation distribution toward a high-probability mean, degrading both performance and diversity (Shumailov et al., 2024; Shimabucoro et al., 2024; Guo et al., 2023). Recent work has pointed towards diversity in training data improving downstream performance, though doesn’t consider the quality of the data in tandem (Chen et al., 2024). In contrast, the BARE pipeline is designed with both of these objectives in mind, producing diverse high-quality data to support model training.
Generation Methods.
Current methods to improve LLM generation diversity include prompting, sampling, and multi-stage generation.
Prompting methods leverage instruction-following by requesting distinct responses in single calls or varying prompts (e.g., assigning distinct ‘personalities’ or randomizing few-shot examples) (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024; Chen et al., 2024; Li et al., 2022). However, compared to BARE, these methods depend on pre-existing data or curated prompts, limiting scalability.
Sampling methods like temperature scaling and nucleus sampling (Holtzman et al., 2020) are widely used, but often prioritize token-level randomness over semantic diversity. Indeed, methods such as logit suppression (Chung et al., 2023) can enhance diversity but may require significant manual refinement to maintain quality.
Multi-Stage generation methods vary in approach. SimpleStrat (Wong et al., 2024) explicitly generates strata of the solution space and then conditions responses on each, effective for short-answer responses (often a single word or phrase), though limiting applicability to broader synthetic data generation. More complex frameworks require extensive human setup or curation to seed the diversity with topics, making them less scalable (Lambert et al., 2024; Li et al., 2024a, 2023). In contrast, BARE, is a practical two-step pipeline that we show works generally with minimal human involvement.
Prior work has studied differences in base and instruct models in calibration (OpenAI, 2024b) and agentic environments (Li et al., 2024b), though no work to our knowledge has leveraged base models for synthetic data generation.
Evaluating Synthetic Data.
Synthetic data utility is typically assessed by downstream performance, though to motivate better systems, we additionally study diversity and entry-wise quality.
Token-level metrics like self-BLEU (Zhu et al., 2018) are common, though embedding-based approaches (e.g., BERTScore (Zhang et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019)) better capture semantic diversity rather than token diversity, which is our primary focus.
Lastly, while dataset-wide quality is often measured via downstream performance, assessing individual synthetic samples remains underexplored. In Section 3, we introduce an entry-wise quality measure to evaluate sample realism, ensuring robust synthetic data generation.
3 Motivation
Synthetic data generation should result in a diverse dataset with high-quality entries. Thus we motivate our design of BARE with investigations on the diversity and entry-wise quality of synthetically generated data from base and instruct-tuned models, independent of downstream utility.
3.1 Diversity & Quality
Diversity.
Following Tevet and Berant (2021), we use the average neural similarity score to measure the diversity of a generated dataset. Specifically, we use OpenAI’s text-embedding-3-small (OpenAI, 2024a) to generate embeddings and use cosine similarity to calculate similarity scores, as recommended by OpenAI. We calculate pairwise cosine similarity scores for items in a generated dataset and analyze the resulting distribution of similarities. A lower average similarity indicates a more diverse dataset.
Entry-wise quality.
We seek a metric measuring the quality of individual entries in the generated datasets. To this end, we propose the indistinguishability rate (IR) of the entries as a quality metric. We believe that we are the first to propose this metric for the quality of an individual data entry. Specifically, we measure how often a strong LLM (e.g. GPT-4o) fails to identify the synthetically generated entry as being of lower quality when combined with $n=3$ real dataset entries. A high IR indicates that generated data closely matches properties of real-world data and the discriminator is randomly guessing, while a low IR indicates the generated data are out-of-distribution from real world data. An example IR prompt is available in Appendix B.
3.2 Experimental Setup
Models.
To investigate diversity and quality differences between instruct-tuned and base models, we evaluate a synthetic dataset generated using Llama-3.1-70B-Instruct and Llama-3.1-70B-Base (Dubey et al., 2024). We additionally use GPT-4o (OpenAI, 2024c) as an instruct-tuned model with strong instruction-following capabilities.
To allow fair comparisons and simulate a low-data setting we always provide three few-shot examples.
Domains.
We examine a variety of domains covering many tasks. Our data generation domains are:
  • Enron Emails (Klimt and Yang, 2004) generating training data for classifying emails as spam or legitimate. We ensure class-balanced synthetic data by explicitly conditioning each generation on a uniform class distribution.
  • 20 Newsgroups (Pedregosa et al., 2011) generating training data for classifying Usenet messages into one of 20 newsgroup sources. The model generates classes along with content, allowing the generation method to determine the class distribution of the synthetic dataset.
  • Retrieval-Augmented Fine-Tuning (RAFT) (Zhang et al., 2024c), a domain and generator-agnostic synthetic data generation framework for fine-tuning data in RAG tasks. Here, Q/A pair data generation needs to be conditioned on contexts mimicking retrieval results. We use:
    • HotpotQA (Yang et al., 2018), a general Wikipedia-based short-answer task.
    • PubMedQA (Jin et al., 2019), a medical abstract-based yes/no/maybe question-answering task.
  • GSM8K (Cobbe et al., 2021), generating grade-school math word problems and solutions for fine-tuning.
  • LiveCodeBench’s Test Output Prediction (LCB TOP) (Jain et al., 2024), generating coding questions and answers on predicting test case outputs given a natural language description of an algorithm’s expected behavior and test input for fine-tuning.
For the classification tasks of Enron and Newsgroups, we generate a dataset of size $n=500$. For the generative model fine-tuning tasks of HotpotQA, PubMedQA, GSM8K, and LCB TOP, we generate a larger dataset of size $n=1000$.
We sample at a default temperature of $0.7$ for base models and the highest temperature at which we experimentally found data generation is still coherent for instruct models, which is $1.0$ for Llama models and $1.2$ for GPT-4o. However, for Enron, we sample from GPT-4o at a temperature of $1.0$ to maintain generation coherence.
Note we do not measure diversity in the RAFT domains as we expect noise due to conditioning on different documents/contexts. For Enron, due to the relatively large differences between spam and legitimate emails, only similarities between emails of the same type are calculated.
3.3 Investigation Results
Figure 3: Distribution of pairwise embedding cosine similarities for Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Enron spam emails. Base model distribution have more density in the low-similarity region and less density in the high-similarity region, indicating greater diversity. Note that frequency is of the cosine similarity score.
Figure 4: Topic coverage of Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Newsgroups. Coverage after $n=500$ generations. The base model covers all topics except one whereas the instruct model covers only $60\%$ of the topics.
Diversity.
Looking at the pairwise cosine similarity distributions of the embeddings of the generated dataset in Figure 3 (and recalling Figure 2), the base distribution is consistently shifted to the left, indicating that the base model generations are more diverse and correspond with a decrease in average similarity (Table 1). The higher diversity of base models is more strikingly shown in the coverage of generated classes for Newsgroups (Figure 4), where over $500$ generations the Instruct model covers only 14 of the 20 classes. In contrast, the Base model covers every class except 1 (talk.religion.misc).
Table 1: Average pairwise embedding cosine similarity of Llama-3.1-70B-Instruct vs. Llama-3.1-70B-Base generated data. Generations from Base are almost always more diverse than Instruct, despite always sampling at a lower temperature ($0.7$ vs $1.0$).
Dataset Instruct Base
Enron 0.450 0.350
Newsgroups 0.256 0.162
GSM8K 0.421 0.313
LCB TOP 0.389 0.468
This trend is reflected in Table 1, with all domains except one showing results for base models with higher diversity. Upon inspection, we attribute the reversal in trend in LCB TOP to the base model repeating phrases from the examples in the prompt. This is related to potential issues with the quality of base model generations, which we discuss below.
Table 2: Average pairwise embedding cosine similarity of various prompting techniques. GPT-4o is used as the prompting generator due to the need for strong instruction following capabilities; Llama models frequently derailed. Llama-3.1-70B-Base generations are almost uniformly more diverse than any prompting method, except for persona prompting on GSM8K, where it is comparable.
Method Enron GSM8K
GPT-4o Independent 0.574 0.427
GPT-4o Persona 0.580 0.308
GPT-4o Sequential 0.511 0.398
GPT-4o In-One 0.363 0.347
GPT-4o Dynamic Fewshot 0.511 0.463
Llama-3.1-70B-Base 0.350 0.313
Sampling Methods
We further explore data generation using several different sampling methods. As base models struggle to follow complex instructions, only instruct models are used for methods beyond independent sampling.
  • Independent (temperature) Sampling: The model generates $n$ samples to form a dataset.
  • Persona Prompting: The model is instructed to respond as a predefined persona (Fröhling et al., 2024).
  • Sequential Prompting: The model iteratively generates outputs distinct from previous ones.
  • In-One Prompting: The model is prompted to generate $k$ different entries in a single response (Zhang et al., 2024b).
  • Dynamic Few-Shot Examples: Few-shot examples are randomly selected for each call (breaking the assumption of low-data conditioned generation) (Li et al., 2022).
Table 2 shows that base models generally yield higher diversity than almost all prompting methods on the average embedding distance metric in two domains. One exception is persona prompting on GSM8K — though this diversity arises more from flavor text differences due to personas rather than actual content.
We thus find that base models are generally more diverse than instruct-tuned models and that temperature increases and prompting methods are insufficient to bridge the gap, motivating our usage of base models in the first stage of BARE.
Figure 5: Indistinguishability Rate for Llama-3.1-70B-Base and Instruct model data generation methods across various datasets. Llama-3.1-70B-Instruct is almost uniformly better at appearing high-quality alongside real-world data.
Entry-wise quality.
Figure 5 presents our results. In general, the instruct-tuned model has a higher IR, indicating that it is better at producing generations that resemble high-quality data. Note that some IRs are well above 75%. This outcome is not unexpected: if a model consistently generates data that aligns with the most common patterns in the real-world distribution, it becomes difficult to distinguish from actual data. Since the real-world entries often also include non-modal (less frequent) samples, a discriminator tasked with identifying lower quality data may instead misclassify these less common real-world samples as synthetic.
As mentioned above, on LCB TOP the base model repeats phrases from examples in the prompt. This leads to a higher IR as the generations copy real-world examples. Thus, while individual base generations are technically more realistic, their shortcomings are captured in the diversity metric.
We therefore find that the superior instruction follow-up capabilities of instruct-tuned models can help generate more realistic data, motivating our usage of instruct-tuned models in the second stage of BARE.
4 Base-Refine (BARE)
Figure 6: Instruct-tuned models provide high-quality but low-diversity data, while base models provide low-quality but high-diversity data. BARE independently generates a diverse initial set of data points with a base model and refines each entry individually with an instruct-tuned model to create a high-quality, high-diversity dataset. In this example of a real grade school math problem generation, the Llama-3.1-70B-Base model hallucinates in its answer to its own question. The refiner (Llama-3.1-70B-Instruct) recognizes this and disambiguates the question and corrects the reasoning.
We leverage our insights to propose BARE, a practical synthetic data generation method combining the diversity of base models with the quality of instruct models. BARE uses a base model to generate an initial set of diverse but potentially lower quality data, after which an instruct-tuned model individually refines each example from the initial set – as shown in the example in Figure 6.
In the base step, the base model uses a prompt with minimal few-shot examples to generate. In the refine step, the instruct-tuned model is instructed to retain the concept of the base model generation while improving it according to specific criteria (e.g., realism, correctness). This retains the overall diversity of the base model generated set while exerting greater control over the quality of the final generation.
Importantly, few-shot examples are only required in the base step to ensure formatting (though they can also be included in the refine step), and only in small amounts. Thus, BARE is especially useful in domains with very limited real-world data as little human effort is required to generate a diverse high-quality synthetic dataset. In our experiments, we limit ourselves to just three few-shot examples. In addition, we intentionally use very general prompts for BARE to demonstrate its flexibility, underscoring the potential for even greater improvement with tailored prompts. Representative prompts are included in Appendix B.
5 Evaluation
We evaluate BARE for diversity, data quality, and downstream utility on the same domains and against the same baseline methods presented in Section 3. We implement BARE with Llama-3.1-70B-Base in the base generation stage and Llama-3.1-70B-Instruct in the refinement stage, and perform additional experiments with the Llama 3.1 8B family and with Llama 3.1 8/70B as the base model and GPT-4o as the refiner (Dubey et al., 2024; OpenAI, 2024c).
We follow generally the same sampling temperatures as in Section 3. However, for Enron and Newsgroups, we instead sample from Llama-3.1-8B-Instruct at temperature $0.7$ in order to maintain coherence in generations. We always sample the refine model at a temperature of $0.7$.
For evaluation of downstream utility for the generative tasks, we LoRA (Hu et al., 2021) fine-tune Llama-3.1-8B-Instruct for 4 epochs using the generated data, except for GSM8K where Llama-3.2-1B-Instruct (Dubey et al., 2024) is fine-tuned instead due to high baseline performance of the 8B model. Other fine-tuning hyperparameters are in Section A.1.1. The fine-tuned models are evaluated on a static $n=100$ test set for HotpotQA and PubMedQA, a static $n=500$ test set for GSM8K, and the full $n=442$ test set for LCB TOP. In this section, we will focus only on the generative tasks; the downstream evaluation process for classification tasks is presented in Section A.1.2 and detailed results for all domains can be found in Section A.2.
Figure 7: Distribution of pairwise embedding similarities for Llama-3.1-70B-Base and Llama 3.1 70B BARE generations for GSM8K (top) and LCB TOP (bottom). The Base and BARE distributions are extremely similar for both tasks, indicating that refinement retains the diversity of the base generations. Note that frequency is of the cosine similarity score.
BARE Quality & Diversity.
We begin by looking at the quality/diversity trends of BARE when compared with previous methods. From the histograms plotted in Figure 7, we can see that BARE effectively does not change the similarity distribution of generated data when compared to the base model at all – it is able to successfully retain the diversity of base generations. Detailed results with average embedding similarity scores can be found in Section A.2.
Figure 8: Indistinguishability Rate for Llama-3.1-70B-Base and BARE data generation across various datasets. BARE consistently improves the quality of data generated from base models.
Simultaneously though, looking at Figure 8, we can see that while retaining diversity, BARE leads to a monotonic increase in the IR for every domain - suggesting that it is able to lift the quality of the generations to be on par or in some cases even surpass directly sampling from an instruct model. Combined, this indicates that BARE is capable of leveraging the diversity of base models and quality of instruct-tuned models in its end generations.
Figure 9: Fine-tuned model accuracy results on HotpotQA, PubMedQA, GSM8K, and LCB TOP for prompting methods using GPT-4o. Prompting methods have mixed effects on downstream performance when compared against standard sampling, while using GPT-4o with simply the Llama-3.1-8B-Base model in BARE improves consistently over all prompting methods.
Figure 10: Accuracy of a Llama-3.1-8B-Instruct Model (HotpotQA, PubMedQA, LCB TOP) and Llama-3.2-1B-Instruct Model (GSM8K) finetuned on synthetic data generated using Base, Instruct, and BARE methods. Note that GSM8K 8B Instruct results are not shown as data generations were derailed.
Fine-tuned Model Accuracy.
We now show the utility of BARE datasets as a whole. In Figure 10, we demonstrate the accuracy of a model fine-tuned on the datasets generated using different methods. Almost uniformly, BARE leads to greater improvements in downstream model quality than generating from either Base or Instruct models, and across all domains BARE-based data leads to the highest fine-tuned model accuracy.
Impressively, in Figure 9, beyond just base and instruct sampling, we show that enhancing GPT-4o with BARE using only a small base model like Llama-3.1-8B-Base produces monotonically better fine-tuning data than all prompting methods discussed in Section 3.3 that use only GPT-4o. These clear results showcase BARE’s ability to out-perform the existing methods for high-quality, diverse data generation.
Focusing on the RAFT domains, BARE improves upon the standard SOTA pipeline for fine-tuning LLMs for RAG by up to 18.4%, as seen with the 8B family on HotpotQA (standard RAFT is implemented in our Instruct generation results). BARE with both model families also outperforms existing RAFT pipelines on PubMedQA. While BARE with the Llama 3.1 70B family does not improve upon RAFT for HotpotQA, switching out the Llama-3.1-70B-Instruct refiner for GPT-4o does lead to an improvement (Section A.2).
On GSM8K, BARE is the only method that provides useful training data. The un-trained model performance was 21.8%, and fine-tuning on BARE generated data achieves accuracies of 24.9% and 32.8% with the Llama 8B and 70B families, respectively. Accuracy when training with data generated by single model methods either decreased accuracy or had little difference. In fact, training on data generated by Llama-3.1-70B-Instruct led to an accuracy of just 17.8%, which BARE with Llama-3.1-70B-Base refined by GPT-4o outperforms by 101% (35.8% accuracy).
[…]
Similar to HotpotQA, on LCB TOP, BARE with the Llama 3.1 70B family shows less of an improvement than the 70B-instruct-only data generation method. Again, using GPT-4o as a refiner instead of Llama-3.1-70B-Instruct gives stronger results. This suggests that with the right refiner choice, BARE consistently provides gains. We show detailed results of refining with GPT-4o in all domains in Section A.2, emphasizing BARE’s consistently strong data generation capabilities.
We also find that temperature ablations do not meaningfully affect the utility of instruct-tuned model generations, especially when compared to gains using BARE. For details, see Section A.3.
Instruct-Instruct Ablation.
A possible challenge is that the gains in performance are due to a multi-step pipeline rather than combining a base and instruct-tuned model. To demonstrate the importance of using a base model in BARE, we performed an ablation where an instruct-tuned model’s generations are refined by an instruct-tuned model on the GSM8K task. The accuracy on the test set after switching Llama-3.1-70B-Base to Instruct in the first step drops from $29.8\%$ to $25.4\%$. This trend held when GPT-4o was the refiner as well, with a drop from $35.8\%$ to $30.8\%$. At the same time, the pattern of little change in the average pairwise similarity before and after refinement remains, indicating that diversity must be introduced in the first stage. Having earlier established that base models are most effective at diversity, we conclude that the use of base models is a necessary component of BARE. Detailed results from this ablation are in Section A.4.
6 Conclusion
6.1 Key Contributions
In summary, in this work we quantitatively investigated the quality and diversity of various synthetic data generation methods. Through this investigation, we find that base models are generally more diverse than instruct-tuned models while instruct-tuned models produce higher quality output than base models, validating hypotheses in the research community. These insights motivate the design of a better system to generate synthetic data, BARE.
Through extensive experiments, we validate the importance of each step in BARE and demonstrate its ability to preserve base model diversity while enhancing output quality. Moreover, by fine-tuning on BARE-generated data for various domains, we underscore BARE’s practical utility, consistently outperforming existing synthetic data generation methods on downstream tasks such as GSM8K and LiveCodeBench, in addition to RAFT, for which we set a new SOTA.
6.2 Future Work
There are a lot of exciting directions for future work. For one, BARE is not necessarily the only way to elicit diversity in the synthetic data generation process. We were able to achieve all of our results with essentially basic prompting, leading us to believe there is room for even more improvement if one was to, for example, fine-tune the refiner specifically for this use case or use stronger refiner models. Introducing additional stages to BARE could also lead to improvements (e.g., additional refinement steps). Beyond multi-stage systems, the design space for diversity is vast – in one direction, training instruct-tuned models with different objective functions that encourage entropy is another area of exploration.
Furthermore, the implications of a lack of high-quality, diverse generations from LLMs go beyond just synthetic data. For example, much of the work in inference time compute relies on models being able to generate sufficiently different possible trajectories and using these diverse trajectories (exploration) with a feedback signal to improve reasoning capabilities (Zhang et al., 2024a; Zelikman et al., 2022), yet only a few methods look explicitly at generator diversity (Wang et al., 2024).
Lastly, the same BARE method demonstrated in this paper for synthetic training data can also be used with no modifications to generate synthetic evaluation sets, which are especially valuable in so many real-world domains with low-data availability.
Acknowledgements
We would like to thank Justin Wong, Peter West, Raj Ammanabrolu, Alex Dimakis, Alex Xu, and Andrew Qin for their many insightful discussions about this work.
Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Microsoft, NVIDIA, Samsung SDS, SAP, and VMware.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Additional Results
A.1 Downstream Evaluation - Additional Details
A.1.1 Fine-Tuning Task Hyperparameters
We list below the fine-tuning hyperparameters that were used in common for HotpotQA, PubMedQA, GSM8K, and LCB TOP. Learning rate was determined independently for each domain via learning rate sweeps (across orders of magnitude); each sweep gave the same optimal learning rate.
  • Learning Rate: 0.001
  • LoRA $\alpha$: 16
  • LoRA Rank: 8
  • LoRA Dropout: 0.0
A.1.2 Classification Task Setup
The generated data is used to train a BERT-based classifier [Devlin et al., 2018] for 2 epochs on Enron and 9 epochs on Newsgroups. The trained models are evaluated on a static test set with $n=500$ examples for each domain.
A.2 Core Experiment Results - All Domains
This appendix contains diversity, IR, and downstream performance results for all core experiments: generation with Llama 3.1 8B and 70B Base and Instruct models, BARE with Llama 3.1 models of both families, and BARE with the use of GPT-4o.
Note that HotpotQA RAFT and PubMedQA RAFT diversity results present here were not presented in Table 1 as we believe the numbers are noisy and not fit for drawing conclusions, due to the use of $100$ different simulated retrieval contexts that generation was conditioned on (as required by RAFT). Not only does this introduce noise to the similarity calculation, but the strong instruction following capability of instruct models allow them to better leverage the inherent diversity in different prompts. However, for completeness, we report the values in the tables in this appendix.
Table 3: Average pairwise embedding cosine similarity, IR, and downstream F1 results on a randomly selected static $n=500$ subset of Enron. A BERT model with a classification head was trained for $2$ epochs on the generated data $(n=500)$. Only pairwise similarities for generations within the same class (spam or legitimate) were calculated.
Generation Method Average Embedding Similarity IR Downstream F1
Llama 3.1 8B Instruct 0.500 86.0% 0.753
Llama 3.1 70B Instruct 0.450 85.0% 0.848
Llama 3.1 8B Base 0.368 63.5% 0.790
Llama 3.1 70B Base 0.350 74.5% 0.819
BARE Llama 3.1 8B 0.413 85.0% 0.872
BARE Llama 3.1 70B 0.406 82.0% 0.771
BARE GPT-4o + Llama 3.1 8B Base 0.379 84.5% 0.872
BARE GPT-4o + Llama 3.1 70B Base 0.356 88.5% 0.846
Table 4: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on on a randomly selected static $n=500$ subset of Newsgroups. A BERT model with a classification head was trained for $9$ epochs on the generated data $(n=500)$.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Llama 3.1 8B Instruct 0.271 85% 26%
Llama 3.1 70B Instruct 0.246 82% 30%
Llama 3.1 8B Base 0.155 58% 41%
Llama 3.1 70B Base 0.162 78% 29%
BARE Llama 3.1 8B 0.162 91% 40%
BARE Llama 3.1 70B 0.134 93% 49%
BARE GPT-4o + Llama 3.1 8B Base 0.131 81% 44%
BARE GPT-4o + Llama 3.1 70B Base 0.285 87% 47%
Table 5: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=100$ subset of HotpotQA RAFT. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 33%
Llama 3.1 8B Instruct 0.214 76% 49%
Llama 3.1 70B Instruct 0.216 90% 55%
Llama 3.1 8B Base 0.221 62% 50%
Llama 3.1 70B Base 0.209 77% 53%
BARE Llama 3.1 8B 0.217 77% 58%
BARE Llama 3.1 70B 0.210 88% 54%
BARE GPT-4o + Llama 3.1 8B Base 0.214 78% 57%
BARE GPT-4o + Llama 3.1 70B Base 0.205 89% 56%
Table 6: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=100$ subset of PubMedQA RAFT. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 53%
Llama 3.1 8B Instruct 0.376 63% 72%
Llama 3.1 70B Instruct 0.373 73% 73%
Llama 3.1 8B Base 0.396 39% 68%
Llama 3.1 70B Base 0.603 31% 73%
BARE Llama 3.1 8B 0.377 47% 73%
BARE Llama 3.1 70B 0.367 72% 79%
BARE GPT-4o + Llama 3.1 8B Base 0.385 57% 72%
BARE GPT-4o + Llama 3.1 70B Base 0.474 40% 63%
Table 7: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a $n=500$ randomly selected static subset of GSM8K. A Llama-3.2-1B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$) instead of the 3.1 8B model due to its high no-training performance. The baseline performance of the Llama-3.2-1B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row. Llama-3.1-8B-Instruct generation results are not reported due to data generation derailing.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 21.8%
Llama 3.1 8B Instruct N/A N/A N/A
Llama 3.1 70B Instruct 0.421 51% 17.8%
Llama 3.1 8B Base 0.310 23% 19.2%
Llama 3.1 70B Base 0.313 44% 22.4%
BARE Llama 3.1 8B 0.310 27% 24.8%
BARE Llama 3.1 70B 0.305 54% 29.8%
BARE GPT-4o + Llama 3.1 8B Base 0.295 60% 32.8%
BARE GPT-4o + Llama 3.1 70B Base 0.302 64% 35.8%
Table 8: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on the full LCB TOP set. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 18.6%
Llama 3.1 8B Instruct 0.416 51% 20.6%
Llama 3.1 70B Instruct 0.389 49% 26.0%
Llama 3.1 8B Base 0.468 36% 24.9%
Llama 3.1 70B Base 0.477 57% 24.4%
BARE Llama 3.1 8B 0.462 47% 28.1%
BARE Llama 3.1 70B 0.481 64% 25.6%
BARE GPT-4o + Llama 3.1 8B Base 0.459 72% 26.7%
BARE GPT-4o + Llama 3.1 70B Base 0.471 68% 27.4%
A.3 Independent Sampling Temperature Ablations - HotpotQA, PubMedQA, and LCB TOP
This appendix contains diversity, IR, and downstream performance results for our temperature ablation experiments. We perform a temperature sweep for Llama-3.1-8B-Instruct generation with $t=0.5,0.7,1.0$. We find that while adjusting the temperature can improve downstream performance, in general the gains are small relative to gains by using BARE.
Table 9: Temperature ablations with independent sampling from Llama-3.1-8B-Instruct. Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=100$ subset of HotpotQA RAFT. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 33%
Llama 3.1 8B Instruct ($t=1.0$) 0.214 76% 49%
Llama 3.1 8B Instruct ($t=0.7$) 0.216 77% 50%
Llama 3.1 8B Instruct ($t=0.5$) 0.220 83% 42%
Llama 3.1 8B Base ($t=1.0$) 0.221 62% 50%
BARE Llama 3.1 8B (All $t=0.7$) 0.217 77% 58%
Table 10: Temperature ablations with independent sampling from Llama-3.1-8B-Instruct. Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=100$ subset of PubMedQA RAFT. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 53%
Llama 3.1 8B Instruct ($t=1.0$) 0.376 62% 72%
Llama 3.1 8B Instruct ($t=0.7$) 0.375 71% 72%
Llama 3.1 8B Instruct ($t=0.5$) 0.377 73% 75%
Llama 3.1 8B Base ($t=1.0$) 0.396 39% 68%
BARE Llama 3.1 8B (All $t=0.7$) 0.377 47% 73%
Table 11: Temperature ablations with independent sampling from Llama-3.1-8B-Instruct. Average pairwise embedding cosine similarity, IR, and downstream accuracy results on the full LCB TOP set. A Llama-3.1-8B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.1-8B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 18.6%
Llama 3.1 8B Instruct ($t=1.0$) 0.365 33% 20.1%
Llama 3.1 8B Instruct ($t=0.7$) 0.416 51% 20.6%
Llama 3.1 8B Instruct ($t=0.5$) 0.450 53% 22.9%
Llama 3.1 8B Base ($t=1.0$) 0.468 36% 24.9%
BARE Llama 3.1 8B (All $t=0.7$) 0.462 47% 28.1%
A.4 BARE First Stage Ablations - GSM8K
This appendix contains diversity, IR, and downstream performance results for our ablation replacing the first stage of BARE with an instruct-tuned model, specifically Llama-3.1-70B-Instruct. We refine using Llama-3.1-70B-Instruct and GPT-4o, and investigate the change in downstream performance compared to standard BARE (using Llama-3.1-70B-Base in the first stage). Note that dataset diversity is unchanged compared to direct generation from Llama-3.1-70B-Instruct, that IR improves after refinement, and that downstream performance is consistently worse than standard BARE.
Table 12: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a $n=500$ randomly selected static subset of GSM8K. A Llama-3.2-1B-Instruct model was fine-tuned for 4 epochs on the generated data ($n=1000$). The baseline performance of the Llama-3.2-1B-Instruct model on the evaluation set prior to any fine-tuning is reported in the first row.
Generation Method Average Embedding Similarity IR Downstream Accuracy
Baseline Performance 21.8%
Llama 3.1 70B Instruct 0.421 51% 22.4%
Llama 3.1 70B Instruct Self-Refine 0.422 63% 25.4%
GPT-4o Refining Llama 3.1 70B Instruct 0.421 70% 30.8%
BARE Llama 3.1 70B 0.305 54% 29.8%
BARE GPT-4o + Llama 3.1 70B Base 0.302 64% 35.8%
Appendix B GSM8K Prompt Examples
In this appendix, we provide exact prompts used for the GSM8K domains, representative of those used throughout this work. Examples are formatted for inclusion in the prompts in the “{examples}” fields, with “EXAMPLE START” and “EXAMPLE END” delimiters for the base prompt. BARE uses the standard Base Prompt in the base generation step.
Static Few-shot Examples
Example 1 Question: Alice has 20 quarters. She wants to exchange them for nickels and so she goes to the bank. After getting back from the bank, she discovers that 20% of the nickels are iron nickels worth \$3 each. What is the total value of her money now? Answer: A quarter is worth five nickels because .25 / .05 = .$\ll.25/.05=5\gg 5≪ .$. She gets 100 nickels from the bank because 20 x 5 = $\ll 20*5=100\gg 100$. 20 of the nickels are iron nickels because 100 x .20 = $\ll 100*.20=20\gg 20$. 80 of the nickels are regular because 100 - 20 = $\ll 100-20=80\gg 80$. The iron nickels are worth \$60 because 20 x 3 = $\ll 20*3=60\gg 60$. The regular nickels are worth \$4 because 80 x .05 = $\ll 80*.05=4\gg 4$. Her money is now worth \$64 because 60 + 4 = $\ll 60+4=64\gg 64$. #### 64
Example 2 Question: A church has 120 members. 40% are adults. The rest are children. How many children more children are there than adults? Answer: There are 48 adults because 120 x .4 = $\ll 120*.4=48\gg 48$. 60% of members are children because 100 - 40 = $\ll 100-40=60\gg 60$. There are 72 children because 120 x .6 = $\ll 120*.6=72\gg 72$. There are 24 more children than adults because 72 - 48 = $\ll 72-48=24\gg 24$. #### 24
Example 3 Question: Lisa is looking to attempt a World Record. She has decided to try and match Joey Chestnut’s record of eating 75 full hotdogs, buns included, in 10 minutes. Halfway through the time Lisa has eaten 20 hotdogs. How many hotdogs will she have to eat per minute to at least tie Joey Chestnut’s record? Answer: Joey Chestnut ate 75 hotdogs to claim the record and Lisa has eaten 20 hot dogs so far, so she still needs to eat 7520=7520=555575-20$\ll$ 75-20=55$\gg$ 5575 - 20 = ≪ 75 - 20 = 55 ≫ 55 hotdogs to tie Joey Chestnut. Lisa has a 10-minute time period to eat the hotdogs and half the time has already passed, which means Lisa has 10/2=10/2=5510/2$\ll$ 10/2=5$\gg$ 510 / 2 = ≪ 10 / 2 = 5 ≫ 5 minutes left until the competition is over. If she needs to eat 55 hotdogs to tie Joey Chestnut and there are 5 minutes left in the competition period, then she needs to eat 55/5=55/5=111155/5$\ll$ 55/5=11$\gg$ 1155 / 5 = ≪ 55 / 5 = 11 ≫ 11 hot dogs per minute to have a chance of tying for a win. #### 11
Base Prompt Here are a few examples of grade school math word problems that require performing a sequence of elementary calculations using basic arithmetic operations. A bright middle school student should be able to solve each problem. The numerical answer is provided at the end of each example after ####. {examples} EXAMPLE START
Instruct Few-shot Prompt Provide an example of a grade school math word problem that requires performing a sequence of elementary calculations using basic arithmetic operations. A bright middle school student should be able to solve each problem. Problems require no concepts beyond the level of early Algebra. You must first specify the question, then provide the very concise reasoning and answer. Provide your example in the following format: Question: [question] Answer: [answer]
Provide only the question and answer in the given format. Note how the numerical answer is provided after #### after each brief reasoning for a question. Here are some examples: {examples} Now it’s your turn. Start your response with the question.
Refine Prompt Improve the given grade school math word problem. Edit the problem or answer to be more similar in style to the examples, and disambiguate as necessary, in addition to correcting any errors. Do not change the theme of the problem. A bright middle school student should be able to solve each problem. Problems require no concepts beyond the level of early Algebra. Note how the numerical answer is provided after #### after each brief reasoning for a question. Provide your edited problem in the following format: Question: [question] Answer: [answer]
Provide only the question and answer in the given format. Here are some examples of categories and problems on those categories: {examples} Now it’s your turn. Here is the question and answer for you to edit: Question: {question} Answer: {answer} Provide only the improved question and answer in the given format. Do not include any commentary or notes. Start your response with the question.
Sequential Prompt Generate a new grade school math word problem that requires performing a sequence of elementary calculations using basic arithmetic operations. A bright middle school student should be able to solve each problem. Problems require no concepts beyond the level of early Algebra. Here are the previously generated examples: {examples} Your new problem should: 1. Be different from the previous examples 2. Follow the same format and style as prior problems Note how the numerical answer is provided after #### after each brief reasoning for a question. Provide only the question and answer in the given format here: Question: [question] Answer: [answer]
Start your response with the question.
In One Prompt Provide {num} examples of problems that might be grade school math word problems that require performing a sequence of elementary calculations using basic arithmetic operations. A bright middle school student should be able to solve each problem. Problems require no concepts beyond the level of early Algebra. You must first specify the question then provide the brief reasoning and answer. Note how the numerical answer is provided after #### after each brief reasoning for a question. Provide your examples in the following format: Question: [question] Answer: [answer]
Here are some examples: {examples} Now it’s your turn. Generate {num} different problems following this format. Your question should be different in content from the examples. Make sure to only provide only the question and answer. Start each example with the question. Delimit the end of an example with the phrase ”END OF EXAMPLE” (all caps) on a new line.
Persona Prompt {persona_description} Provide an example of a grade school math word problem that requires performing a sequence of elementary calculations using basic arithmetic operations. A bright middle school student should be able to solve each problem. Problems require no concepts beyond the level of early Algebra. You must first specify the question, then provide the very concise reasoning and answer. Provide your example in the following format: Question: [question] Answer: [answer]
Provide only the question and answer in the given format. Note how the numerical answer is provided after #### after each brief reasoning for a question. Here are some examples: {examples} Now it’s your turn. Start your response with the question.
Indistinguishability Rate Prompt System Prompt: You are an expert at evaluating question and answer pairs for grade school math word problems. You will be shown {$k$} examples. Each example consists of some context, a question, and an answer. All but one of them is generated from a high quality AI while one of them is of low quality. Your task is to identify which example (1, 2, …, {$k$}) appears to be of low quality. Consider factors like: 1. Differing natural language patterns and flow 2. Differing question structure, clarity, and difficulty 3. Context and specificity 4. Any subtle artifacts or unnatural patterns Analyze each example carefully and explain your reasoning. End with ’Answer: [Question Number]’ where Question Number is 1, 2, …, {$k$}. User Prompt: Here are {$k$} examples. One of them is of low quality. Please identify which one: {questions} Analyze each example and explain which one you think is of low quality. End with ’Answer: [Question Number]’.

### v3

@@
-BARE: Combining Base and Instruction-Tuned Language Models for Better Synthetic Data Generation
+BARE: Leveraging Base Language Models for Few-Shot Synthetic Data Generation
 Abstract
 As the demand for high-quality data in model training grows, researchers and developers are increasingly generating synthetic data to tune and train LLMs.
-A common assumption about synthetic data is that sampling from instruct-tuned models is sufficient; however, these models struggle to produce diverse outputs—a key requirement for generalization.
-Despite various prompting methods, in this work we show that achieving meaningful diversity from instruct-tuned models remains challenging.
-In contrast, we find base models without post-training exhibit greater diversity, but are less capable at instruction following and hence of lower quality.
-Leveraging this insight, we propose Base-Refine (BARE), a synthetic data generation method that combines the diversity of base models with the quality of instruct-tuned models through a two-stage process.
-With minimal few-shot examples and curation, BARE generates diverse and high-quality datasets, improving downstream task performance.
-We show that fine-tuning with as few as 1,000 BARE-generated samples can reach performance comparable to the best similarly sized models on LiveCodeBench tasks.
-Furthermore, fine-tuning with BARE-generated data achieves a 101% improvement over instruct-only data on GSM8K and a 18.4% improvement over SOTA methods on RAFT.
-Machine Learning, ICML
-Figure 1: Histogram of pairwise embedding similarities for 1000 Llama-3.1-70B-Base vs Instruct generations of grade school math problems.
-The base distribution is further to the left, indicating lower similarity and hence higher diversity.
-Note that frequency is of the cosine similarity score.
-Figure 2: Accuracy of a Llama-3.2-1B-Instruct model fine-tuned on 4 different sets of synthetically generated math problems, evaluated on a randomly selected $n=500$ subset of GSM8K.
-Training with BARE-generated data outperforms both Base and Instruct generations.
+However, current data generation methods rely on seed sets containing tens of thousands of examples to prompt instruction-tuned models.
+This reliance can be especially problematic when the curation of high-quality examples is expensive or difficult.
+In this paper we explore the novel few-shot synthetic data generation setting – generating a high-quality dataset from a few examples.
+We show that when working with only a few seed examples, instruction-tuned models used in current synthetic data methods produce insufficient diversity for downstream tasks.
+In contrast, we show that base models without post-training, largely untapped for synthetic data generation, offer substantially greater output diversity, albeit with lower instruction following abilities.
+Leveraging this insight, we propose Base-Refine (BARE), a novel two-stage method that combines the diversity of base models with the quality assurance of instruction-tuned models.
+BARE excels in few-shot synthetic data generation: using only 3 seed examples it generates diverse, high-quality datasets that significantly improve downstream task performance.
+We show that fine-tuning Llama 3.1 8B with 1,000 BARE-generated samples achieves performance comparable to state-of-the-art similarly sized models on LiveCodeBench tasks.
+Furthermore, data generated with BARE enables a 101% improvement for a fine-tuned Llama 3.2 1B on GSM8K over data generated by only instruction-models, and an 18.4% improvement for a fine-tuned Llama 3.1 8B over the state-of-the-art RAFT method for RAG data generation.
 1 Introduction
 As Large Language Models (LLMs) grow in size and capability, the demand for high-quality, diverse data in model training is outpacing human-generated data, necessitating the use of synthetically generated data (Villalobos et al., 2024).
-Consequently, we expect the compute spent on curating and generating data for model training to increase significantly in the coming years, especially in low-data domains where high-quality data is scarce.
-For example, it is common to use LLMs to generate synthetic data for a variety of tasks such as math, code, function calling, general reasoning, etc.
-(Yu et al., 2024; Guo et al., 2024; Patil et al., 2023; Liu et al., 2024).
-The ease of data generation has led many dataset creators (Samvelyan et al., 2024; Zhang et al., 2024c; NovaSky Team, 2025) and model creators (Dubey et al., 2024; Qwen et al., 2025; Nvidia et al., 2024; Guan et al., 2025; NovaSky Team, 2025) to turn to synthetic data sampled from instruct-tuned language models as a replacement for tasks where human-generated data is lacking.
-Challenges with Instruct-Tuned Models.
-Though instruct-tuned models have become more and more capable at complex tasks resulting in higher quality generations, there have been discussions of how accepted methods for post-training can lead to mode collapse (Shumailov et al., 2024; Wong et al., 2024; Lambert et al., 2024).
-Mode collapse refers to an inability for models to generate diverse outputs to queries that don’t have a single response, yet synthetic data is primarily useful when it is both high quality and diverse (Chen et al., 2024; Raventós et al., 2023).
-Intuitively, one might expect a variety of prompting techniques that explicitly encourage diversity to address this gap.
-For example, prompting methods we baseline against include conditioning a generation on all previous ones and prompting for it to be distinct, assuming different personas for different generations, or asking for multiple distinct entries in a single call (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024).
-However, we find that the amount of meaningful diversity from such methods is limited and cannot fully address the gap introduced by post-training.
-On an open-ended generation task like email generation, such prompting methods do not meaningfully improve dataset diversity (Table 2) or trained model accuracy (Figure 9).
-Base vs. Instruct-Tuned Models.
-An alternative approach is to leverage base models, which are not constrained by post-training biases and thus produce outputs that better represent diversity present in real-world data (OpenAI, 2024b).
-More quantitatively, as shown in Figure 2, base models generate outputs with noticeably lower pairwise cosine similarity (mean similarity: $0.313$) compared to instruct-tuned models (mean similarity: $0.421$).
-Here we use cosine similarity as a quantitative surrogate for diversity — a dataset with greater cosine similarity has more similar items and therefore less diversity.
-The improved diversity in base model generations can improve downstream task performance, with Figure 2 showing that the accuracy of Llama-3.2-1B-Instruct on GSM8K increases to 22.5% from 17.8% when fine-tuning on the Llama-3.1-70B-Base generated dataset of word problems rather than the Llama-3.1-70B-Instruct generated dataset.
+Already, it is common to use LLMs to generate synthetic data for a variety of tasks such as math, code, function calling, and general reasoning (Yu et al., 2024; Guo et al., 2024; Patil et al., 2023; Liu et al., 2024).
+The convenience of data generation has led many dataset creators (Samvelyan et al., 2024; Zhang et al., 2024b; NovaSky Team, 2025) and model creators (Dubey et al., 2024; Qwen, 2025; Nvidia, 2024; Guan et al., 2025; NovaSky Team, 2025) to turn to synthetic data sampled from instruction-tuned language models as a replacement for tasks where human-generated data is lacking.
+Figure 1: Histogram of pairwise embedding cosine similarity scores for 1000 Llama-3.1-70B-Base vs Instruct generations of grade school math problems with 3 seed examples.
+The base distribution is further left, indicating lower similarity and hence higher diversity.
+Figure 2: Accuracy of a Llama-3.2-1B-Instruct model fine-tuned on real-world data and 5 different sets of math problems synthetically generated using 3 seed examples, evaluated on a randomly selected $n=500$ subset of GSM8K.
+Training with BARE-generated data outperforms all other data sources.
+Synthetic data is most valuable when it is both high quality and diverse (Chen et al., 2024; Raventós et al., 2023).
+However, while instruction-tuned models excel at generating high-quality outputs, they struggle with generating diverse outputs.
+Recent work suggests that this lack of diversity is a consequence of post-training, where standard techniques lead to mode collapse (Shumailov et al., 2024; Wong et al., 2024; Lambert et al., 2024), limiting a model’s ability to generate varied responses to open-ended queries.
+To achieve sufficient diversity with instruction-tuned models, prevailing methods require developers to provide large and diverse seed sets.
+For instance, OSS-Instruct (Wei et al., 2024) uses 80,000 code snippets and Humpback (Li et al., 2024c) over 500,000 text segments.
+However, this reliance imposes significant burdens on developers through the need for data curation, ultimately hindering adoption.
+While this inefficiency is a general challenge, it becomes especially acute where large seed sets are difficult to obtain (e.g., specialized tasks such as clinical research (Sivarajkumar et al., 2023)).
+Additionally, we show that methods to induce diversity via prompting methods (e.g.
+including past examples in context) (Zhang et al., 2024a; Naik et al., 2023; Fröhling et al., 2024) are insufficient for few-shot generation.
+Thus, a critical need exists for methods that can generate high-quality, diverse data in few-shot settings (i.e., with minimal seed input).
+Base Models as an Alternative.
+Given the limitations of instruction-tuned models in few-shot scenarios, an underexplored alternative source of diversity is the use of base models.
+Not being subject to the same post-training procedures, base models may produce outputs that better reflect the diversity found in real-world data OpenAI (2024b).
+Quantitatively, base models demonstrate greater diversity.
+As shown in Figure 2, they generate outputs with noticeably lower pairwise embedding cosine similarity (mean 0.313) compared to instruction-tuned models (mean 0.421); lower cosine similarity indicates greater diversity.
+This increased diversity can benefit downstream tasks: Figure 2 shows that fine-tuning Llama-3.2-1B-Instruct on GSM8K using data from Llama-3.1-70B-Base yields a 22.5% accuracy, compared to 17.8% using data from Llama-3.1-70B-Instruct.
+However, this advantage comes with a trade-off.
+Due to their weaker instruction-following capabilities, base model generations often suffer from lower quality (Figure 4), which can negate the benefits of increased diversity and hinder downstream training performance.
+Our key insight is that by managing the quality of individual data points using instruction-tuned models, we can harness the diversity advantage of base models effectively.
 BARE.
-In this work, we observe that diversity is critical for synthetic data generation when used in downstream training and fine-tuning tasks.
-We find that while the data generated from a base model may be more diverse, it also tends to be of lower quality (Figure 5).
-This can counteract the gains from diversity and hinder downstream training.
-Our key insight is that by combining base and instruct-tuned models, one can improve the diversity of a dataset while controlling the quality of individual data entries.
-To this end, we introduce Base-Refine (BARE) — a novel approach that leverages the diverse data generated by the base models and refines it with instruct-tuned models.
-In a variety of settings, we find this two-stage process enhances diversity without compromising quality, enabling the generation of datasets that improve downstream performance.
-We evaluate fine-tuning with BARE-generated data on contemporary tasks, such as the recently introduced Retrieval-Augmented Fine-Tuning (RAFT) method (Zhang et al., 2024c).
-RAFT generates synthetic question/answer pairs starting from just a corpus to fine-tune a RAG model, but we find that the diversity of data it generates is low.
-By replacing the data generator in RAFT with BARE, we show we can achieve up to a 18.4% fine-tuned accuracy lift over the existing work that solely uses instruct-tuned models.
-As another example, BARE improves the generation of synthetic math training data.
-We generated grade school math problems similar to GSM8K problems with several methods, shown in Figure 2.
-We find that refining Llama-3.1-70B-Base data generations with Llama-3.1-70B-Instruct improves a fine-tuned model’s accuracy on GSM8K from 22.4% (base) to 29.8% (refined).
-Likewise, using GPT-4o as a generator out of the box (22.4%) does not perform as well as using GPT-4o as a refiner of generations from Llama-3.1-70B-Base (35.8%).
+Leveraging the distinct strengths of base and instruction-tuned models, we introduce Base-Refine (BARE): a novel approach for few-shot synthetic data generation.
+BARE leverages a base model to generate diverse initial generations with minimal seed data and refines each initial generation according to desiderata with instruction-tuned models.
+In a variety of few-shot settings, we show our two-stage process enhances diversity without compromising quality, enabling the generation of datasets that improve downstream performance using only 3 seed examples.
+On the LiveCodeBench Test Output Prediction task, fine-tuning with 1,000 BARE-generated examples achieves performance comparable to state-of-the-art models of similar size.
+Further, when applied to the Retrieval-Augmented Fine-Tuning (RAFT) method (Zhang et al., 2024b)—a state-of-the-art technique for generating synthetic Q&A pairs over retrieved documents for RAG—replacing RAFT’s instruction-tuned only generator with BARE improves the fine-tuned model’s accuracy by up to 18.4% over the original RAFT approach (Figure 8).
+In mathematical reasoning, as exemplified by synthetic GSM8K-style problems, BARE also demonstrates considerable benefits.
+Using Llama-3.1-70B-Base for initial generation and GPT-4o for refinement with BARE increases a fine-tuned Llama-3.2-1B-Instruct model’s accuracy on GSM8K to 35.8%.
+Training on BARE-generated data significantly surpasses the 22.4% accuracy achieved with instruct-only generation using GPT-4o.
+Furthermore, it exceeds the 26.6% accuracy obtained by training on examples from the official human-generated GSM8K training set (Figure 2).
 To summarize, our contributions are:
   1. 1.
-    We quantitatively investigate the quality and diversity of base and instruct-tuned models for synthetic data generation across various sampling methods to motivate better system design.
-In Section 3, we find that base models tend to produce more diverse responses whereas instruct-tuned models offer higher quality.
+    We quantitatively investigate the quality and diversity of base and instruction-tuned models for synthetic data generation across various sampling methods to motivate better system design.
+We show that base models tend to produce far more diverse responses whereas instruction-tuned models offer higher quality.
   2. 2.
-    Using these insights, in Section 4, we propose Base-Refine (BARE), a practical new method for generating synthetic data.
-We show in Section 5 that BARE consistently improves fine-tuned model performance over various baselines, including SOTA generation methods, and with only 1000 samples can lead to fine-tuned models comparable to SOTA models of similar sizes.
+    Building on these insights, we propose Base-Refine (BARE), a practical method novelly leveraging base models for the under-explored task of few-shot synthetic data generation.
+We demonstrate that BARE consistently outperforms past methods on small seed sets, including SOTA generation methods.
+Generating from as few as 3 seed examples, BARE can yield fine-tuned models comparable to SOTA models of similar sizes trained with much larger seed sets.
 2 Related Work
 Synthetic Data.
-Synthetic data is commonly used to train state of the art models like Llama 3.3 (Dubey et al., 2024), Qwen 2.5 (Qwen et al., 2025), Nemotron 4 (Nvidia et al., 2024), and o3-mini (Guan et al., 2025).
+Synthetic data is commonly used to train state of the art models like Llama 3.3 (Dubey et al., 2024), Qwen 2.5 (Qwen, 2025), Nemotron 4 (Nvidia, 2024), and o3-mini (Guan et al., 2025).
 However, prior work has shown its usage poses risks such as model collapse, where iterative training on low-diversity data shifts the generation distribution toward a high-probability mean, degrading both performance and diversity (Shumailov et al., 2024; Shimabucoro et al., 2024; Guo et al., 2023).
-Recent work has pointed towards diversity in training data improving downstream performance, though doesn’t consider the quality of the data in tandem (Chen et al., 2024).
-In contrast, the BARE pipeline is designed with both of these objectives in mind, producing diverse high-quality data to support model training.
+Recent work indicates that diversity in training data can improve downstream performance (Chen et al., 2024).
+However, this research often does not simultaneously consider the quality of the data in tandem.
+In contrast, the BARE pipeline is designed with the twin objectives of diversity and quality in mind, producing diverse, high-quality data to support model training.
 Generation Methods.
-Current methods to improve LLM generation diversity include prompting, sampling, and multi-stage generation.
-Prompting methods leverage instruction-following by requesting distinct responses in single calls or varying prompts (e.g., assigning distinct ‘personalities’ or randomizing few-shot examples) (Zhang et al., 2024b; Naik et al., 2023; Fröhling et al., 2024; Chen et al., 2024; Li et al., 2022).
-However, compared to BARE, these methods depend on pre-existing data or curated prompts, limiting scalability.
-Sampling methods like temperature scaling and nucleus sampling (Holtzman et al., 2020) are widely used, but often prioritize token-level randomness over semantic diversity.
-Indeed, methods such as logit suppression (Chung et al., 2023) can enhance diversity but may require significant manual refinement to maintain quality.
-Multi-Stage generation methods vary in approach.
-SimpleStrat (Wong et al., 2024) explicitly generates strata of the solution space and then conditions responses on each, effective for short-answer responses (often a single word or phrase), though limiting applicability to broader synthetic data generation.
-More complex frameworks require extensive human setup or curation to seed the diversity with topics, making them less scalable (Lambert et al., 2024; Li et al., 2024a, 2023).
-In contrast, BARE, is a practical two-step pipeline that we show works generally with minimal human involvement.
-Prior work has studied differences in base and instruct models in calibration (OpenAI, 2024b) and agentic environments (Li et al., 2024b), though no work to our knowledge has leveraged base models for synthetic data generation.
+Sampling methods like temperature scaling and nucleus sampling (Holtzman et al., 2020) are widely used to improve diversity, but often prioritize token-level randomness over semantic diversity.
+Indeed, methods such as logit suppression (Chung et al., 2023) can enhance diversity but may require significant manual refinement to maintain quality unlike BARE which maintains quality as a first order requirement.
+Many current synthetic data generation processes instead rely on varying the prompt to elicit a diverse dataset.
+One common approach is to rely on heavily curated prompts limiting scalability (Zhang et al., 2024a; Naik et al., 2023; Fröhling et al., 2024; Chen et al., 2024; Li et al., 2022).
+Another effective approach is to utilize a large diverse seed set, up to 100,000s of examples in size (Lambert et al., 2024; Li et al., 2023; Wei et al., 2024; Li et al., 2024c), requiring significant curation effort.
+BARE obviates both of these concerns by focusing on the few-shot setting (we use only 3 seed examples) and showing significant performance gains without prompt engineering.
+Others have explored using an LLM to generate the seed set (Wong et al., 2024; Li et al., 2024a), but still use instruction-tuned models.
+These methods thus still suffer from the limited diversity of instruction-tuned models.
+In contrast, BARE utilizes base models to generate a large and diverse seed set.
+To our knowledge, no prior work has focused on this approach.
+Though prior work has studied differences in base and instruct models in calibration (OpenAI, 2024b) and agentic environments (Li et al., 2024b), no work has leveraged base models for synthetic data.
 Evaluating Synthetic Data.
-Synthetic data utility is typically assessed by downstream performance, though to motivate better systems, we additionally study diversity and entry-wise quality.
-Token-level metrics like self-BLEU (Zhu et al., 2018) are common, though embedding-based approaches (e.g., BERTScore (Zhang et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019)) better capture semantic diversity rather than token diversity, which is our primary focus.
+The utility of synthetic data is typically assessed by downstream performance.
+However, to motivate the development of better systems, we also study diversity and the quality of individual entries (entry-wise quality).
+Token-level metrics like self-BLEU (Zhu et al., 2018) are common, though embedding-based approaches that are commonly used (e.g., BERTScore (Zhang et al., 2019), Sentence-BERT (Reimers and Gurevych, 2019)) better capture semantic diversity rather than token diversity, which is our primary focus.
 Lastly, while dataset-wide quality is often measured via downstream performance, assessing individual synthetic samples remains underexplored.
@@
 3 Motivation
-Synthetic data generation should result in a diverse dataset with high-quality entries.
-Thus we motivate our design of BARE with investigations on the diversity and entry-wise quality of synthetically generated data from base and instruct-tuned models, independent of downstream utility.
-3.1 Diversity & Quality
+Synthetic data generation should result in a diverse and high-quality dataset.
+In this section, we run experiments to demonstrate the challenges of few-shot synthetic data generation with current methods and the differences between base and instruction tuned models.
+We demonstrate the potential of base models as a source of diversity, motivating the design of BARE.
+3.1 Diversity & Quality Metrics
 Diversity.
-Following Tevet and Berant (2021), we use the average neural similarity score to measure the diversity of a generated dataset.
+Following Tevet and Berant (2021); Cann et al.
+(2023); Cox et al.
+(2021), we use the average neural similarity score to measure the diversity of a generated dataset.
 Specifically, we use OpenAI’s text-embedding-3-small (OpenAI, 2024a) to generate embeddings and use cosine similarity to calculate similarity scores, as recommended by OpenAI.
@@
 Entry-wise quality.
-We seek a metric measuring the quality of individual entries in the generated datasets.
-To this end, we propose the indistinguishability rate (IR) of the entries as a quality metric.
-We believe that we are the first to propose this metric for the quality of an individual data entry.
-Specifically, we measure how often a strong LLM (e.g.
-GPT-4o) fails to identify the synthetically generated entry as being of lower quality when combined with $n=3$ real dataset entries.
-A high IR indicates that generated data closely matches properties of real-world data and the discriminator is randomly guessing, while a low IR indicates the generated data are out-of-distribution from real world data.
-An example IR prompt is available in Appendix B.
+To measure entry-wise quality, we propose the indistinguishability rate (IR), a novel metric inspired by the adversarial framework of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014).
+We task a strong LLM (e.g., GPT-4o) as a "discriminator" to distinguish a synthetic entry from $n=3$ real entries.
+The IR is the rate at which it fails to distinguish the synthetic entry.
+A high IR suggests the synthetic data closely mimics real data, while a low IR indicates it is easily identifiable as out-of-distribution.
+An example IR prompt is in Appendix D.
 3.2 Experimental Setup
 Models.
-To investigate diversity and quality differences between instruct-tuned and base models, we evaluate a synthetic dataset generated using Llama-3.1-70B-Instruct and Llama-3.1-70B-Base (Dubey et al., 2024).
-We additionally use GPT-4o (OpenAI, 2024c) as an instruct-tuned model with strong instruction-following capabilities.
-To allow fair comparisons and simulate a low-data setting we always provide three few-shot examples.
+To investigate diversity and quality differences between instruction-tuned and base models, we evaluate synthetic datasets generated using Llama-3.1-70B-Instruct and Llama-3.1-70B-Base (Dubey et al., 2024).
+We also use GPT-4o (OpenAI, 2024c) as a stronger instruction-tuned model.
 Domains.
-We examine a variety of domains covering many tasks. Our data generation domains are:
+To analyze the few-shot setting, each generation method is exposed to only three real-world examples from the following established benchmarks:
   • Enron Emails (Klimt and Yang, 2004) generating training data for classifying emails as spam or legitimate.
@@
   • 20 Newsgroups (Pedregosa et al., 2011) generating training data for classifying Usenet messages into one of 20 newsgroup sources.
-The model generates classes along with content, allowing the generation method to determine the class distribution of the synthetic dataset.
-  • Retrieval-Augmented Fine-Tuning (RAFT) (Zhang et al., 2024c), a domain and generator-agnostic synthetic data generation framework for fine-tuning data in RAG tasks.
-Here, Q/A pair data generation needs to be conditioned on contexts mimicking retrieval results.
+We generate classes along with content, allowing the generation method to determine the class distribution of the synthetic dataset.
+  • Retrieval-Augmented Fine-Tuning (RAFT) (Zhang et al., 2024b), a domain and generator-agnostic synthetic data generation framework for fine-tuning data in RAG tasks.
+Q/A pair generation is conditioned on contexts mimicking retrieval results from a corpus.
 We use:
@@
     • PubMedQA (Jin et al., 2019), a medical abstract-based yes/no/maybe question-answering task.
-  • GSM8K (Cobbe et al., 2021), generating grade-school math word problems and solutions for fine-tuning.
+  • GSM8K (Cobbe et al., 2021), generating grade-school math problems and solutions for fine-tuning.
   • LiveCodeBench’s Test Output Prediction (LCB TOP) (Jain et al., 2024), generating coding questions and answers on predicting test case outputs given a natural language description of an algorithm’s expected behavior and test input for fine-tuning.
@@
 For the generative model fine-tuning tasks of HotpotQA, PubMedQA, GSM8K, and LCB TOP, we generate a larger dataset of size $n=1000$.
-We sample at a default temperature of $0.7$ for base models and the highest temperature at which we experimentally found data generation is still coherent for instruct models, which is $1.0$ for Llama models and $1.2$ for GPT-4o.
-However, for Enron, we sample from GPT-4o at a temperature of $1.0$ to maintain generation coherence.
-Note we do not measure diversity in the RAFT domains as we expect noise due to conditioning on different documents/contexts.
-For Enron, due to the relatively large differences between spam and legitimate emails, only similarities between emails of the same type are calculated.
+For the RAFT domains, this would mean 10 questions for each of 100 simulated retrievals; we do not compute diversity for RAFT due to differences in retrievals.
+We use straight-forward prompts with limited domain-tailoring (examples in Appendix D).
+Note that our prompt for instruction-tuned models is comparable to those used in current synthetic data generation techniques such as OSS-Instruct (Wei et al., 2024).
+Indeed, validation training runs adapting OSS-Instruct’s prompts to our setting yield comparable results to our prompts.
+We thus use our instruction-tuned-only methods as a stand-in for existing methods.
+Further sampling details, such as temperature, are in Appendix A.
+Figure 3: Distribution of pairwise embedding cosine similarity scores for Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Enron spam.
+Base model distribution have more density in the low-similarity region and less density in the high-similarity region, indicating greater diversity.
 3.3 Investigation Results
-Figure 3: Distribution of pairwise embedding cosine similarities for Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Enron spam emails.
-Base model distribution have more density in the low-similarity region and less density in the high-similarity region, indicating greater diversity.
-Note that frequency is of the cosine similarity score.
-Figure 4: Topic coverage of Llama-3.1-70B-Instruct and Llama-3.1-70B-Base generations for Newsgroups.
-Coverage after $n=500$ generations.
-The base model covers all topics except one whereas the instruct model covers only $60\%$ of the topics.
 Diversity.
-Looking at the pairwise cosine similarity distributions of the embeddings of the generated dataset in Figure 3 (and recalling Figure 2), the base distribution is consistently shifted to the left, indicating that the base model generations are more diverse and correspond with a decrease in average similarity (Table 1).
-The higher diversity of base models is more strikingly shown in the coverage of generated classes for Newsgroups (Figure 4), where over $500$ generations the Instruct model covers only 14 of the 20 classes.
-In contrast, the Base model covers every class except 1 (talk.religion.misc).
+From the pairwise cosine similarity distributions of the embeddings in Figure 3 (and recalling Figure 2), we see the base distribution is consistently further to the left, indicating that the base model generations are more diverse.
+The better diversity of base models is further reflected in Table 1, with all domains except one showing lower mean similarity for base models, indicating higher diversity.
+Upon inspection, we attribute the reversal in trend in LCB TOP to the base model repeating phrases from the examples.
+This is related to potential issues with the quality of base model generations, which we discuss below.
 Table 1: Average pairwise embedding cosine similarity of Llama-3.1-70B-Instruct vs.
@@
 Generations from Base are almost always more diverse than Instruct, despite always sampling at a lower temperature ($0.7$ vs $1.0$).
-Dataset Instruct Base
-Enron 0.450 0.350
-Newsgroups 0.256 0.162
-GSM8K 0.421 0.313
-LCB TOP 0.389 0.468
-This trend is reflected in Table 1, with all domains except one showing results for base models with higher diversity.
-Upon inspection, we attribute the reversal in trend in LCB TOP to the base model repeating phrases from the examples in the prompt.
-This is related to potential issues with the quality of base model generations, which we discuss below.
+Enron Newsgroups GSM8K LCB TOP
+Instruct 0.450 0.256 0.421 0.389
+Base 0.350 0.162 0.313 0.468
+In addition to repeated independent sampling from the instruction-tuned model, we further explore diversity-eliciting prompting for data generation.
+We do not explore such methods on base models due to their poor instruction-following capabilities.
+The methods we examine are:
+  • Persona Prompting: Model responds as a predefined persona (Fröhling et al., 2024).
+  • Sequential Prompting: Model iteratively generates outputs distinct from previous ones.
+  • In-One Prompting: Model generates $k$ different entries in a single response (Zhang et al., 2024a).
+  • Dynamic Few-Shot Examples: Different few-shot examples are randomly selected for each call (using a larger seed set) (Li et al., 2022).
 Table 2: Average pairwise embedding cosine similarity of various prompting techniques.
@@
 Llama-3.1-70B-Base generations are almost uniformly more diverse than any prompting method, except for persona prompting on GSM8K, where it is comparable.
-Method Enron GSM8K
-GPT-4o Independent 0.574 0.427
-GPT-4o Persona 0.580 0.308
-GPT-4o Sequential 0.511 0.398
-GPT-4o In-One 0.363 0.347
-GPT-4o Dynamic Fewshot 0.511 0.463
-Llama-3.1-70B-Base 0.350 0.313
-Sampling Methods
-We further explore data generation using several different sampling methods.
-As base models struggle to follow complex instructions, only instruct models are used for methods beyond independent sampling.
-  • Independent (temperature) Sampling: The model generates $n$ samples to form a dataset.
-  • Persona Prompting: The model is instructed to respond as a predefined persona (Fröhling et al., 2024).
-  • Sequential Prompting: The model iteratively generates outputs distinct from previous ones.
-  • In-One Prompting: The model is prompted to generate $k$ different entries in a single response (Zhang et al., 2024b).
-  • Dynamic Few-Shot Examples: Few-shot examples are randomly selected for each call (breaking the assumption of low-data conditioned generation) (Li et al., 2022).
-Table 2 shows that base models generally yield higher diversity than almost all prompting methods on the average embedding distance metric in two domains.
-One exception is persona prompting on GSM8K — though this diversity arises more from flavor text differences due to personas rather than actual content.
-We thus find that base models are generally more diverse than instruct-tuned models and that temperature increases and prompting methods are insufficient to bridge the gap, motivating our usage of base models in the first stage of BARE.
-Figure 5: Indistinguishability Rate for Llama-3.1-70B-Base and Instruct model data generation methods across various datasets.
-Llama-3.1-70B-Instruct is almost uniformly better at appearing high-quality alongside real-world data.
+Prompting GPT-4o Llama-3.1-70B-Base
+Ind. Persona Seq. In-One Dyn. Fewshot Independent
+Enron 0.574 0.580 0.511 0.363 0.511 0.350
+GSM8K 0.427 0.308 0.398 0.347 0.463 0.313
+Table 2 shows that base models generally yield higher diversity than almost all prompting methods in the two domains, as measured by the average embedding distance.
+The exception is persona prompting on GSM8K — though this diversity arises more from flavor text differences due to personas rather than actual content.
+We thus find that base models are generally more diverse than instruction-tuned models and that temperature increases and prompting methods are insufficient to bridge the gap, motivating our usage of base models in the first stage of BARE.
+Figure 4: Indistinguishability Rate for Llama-3.1-70B Base and Instruct data generation methods across various datasets.
+Llama-3.1-70B-Instruct is almost uniformly better at generating examples that appear in-domain when compared alongside real-world data.
 Entry-wise quality.
 Figure 4 presents our results.
-In general, the instruct-tuned model has a higher IR, indicating that it is better at producing generations that resemble high-quality data.
-Note that some IRs are well above 75%.
-This outcome is not unexpected: if a model consistently generates data that aligns with the most common patterns in the real-world distribution, it becomes difficult to distinguish from actual data.
+In general, the instruction-tuned model has a higher indistinguishability rate (IR) (see Section 3.1), indicating that it is better at producing generations that resemble high-quality data.
+Some IRs are well above 75%, which is not unexpected: if a model consistently generates data that aligns with the most common patterns in the real-world distribution, it becomes difficult to distinguish from actual data.
 Since the real-world entries often also include non-modal (less frequent) samples, a discriminator tasked with identifying lower quality data may instead misclassify these less common real-world samples as synthetic.
-As mentioned above, on LCB TOP the base model repeats phrases from examples in the prompt.
-This leads to a higher IR as the generations copy real-world examples.
-Thus, while individual base generations are technically more realistic, their shortcomings are captured in the diversity metric.
-We therefore find that the superior instruction follow-up capabilities of instruct-tuned models can help generate more realistic data, motivating our usage of instruct-tuned models in the second stage of BARE.
+As mentioned above, on LCB TOP the base model repeats phrases from examples in the prompt, leading to a higher IR as the repeated examples are real.
+Consequently, while individual base generations are technically more realistic, their shortcomings are captured in the diversity metric.
+We thus find that the superior instruction following capabilities of instruction-tuned models generates more realistic data, motivating our usage of instruction-tuned models in the second stage of BARE.
 4 Base-Refine (BARE)
-Figure 6: Instruct-tuned models provide high-quality but low-diversity data, while base models provide low-quality but high-diversity data.
-BARE independently generates a diverse initial set of data points with a base model and refines each entry individually with an instruct-tuned model to create a high-quality, high-diversity dataset.
+Figure 5: BARE combines base models with instruction tuned models.
+Instruction-tuned models provide high-quality but low-diversity data, while base models provide low-quality but high-diversity data.
+With minimal seed examples, BARE independently generates a diverse initial set of data points with a base model and refines each entry individually with an instruction-tuned model to create a high-quality, high-diversity dataset.
 In this example of a real grade school math problem generation, the Llama-3.1-70B-Base model hallucinates in its answer to its own question.
 The refiner (Llama-3.1-70B-Instruct) recognizes this and disambiguates the question and corrects the reasoning.
-We leverage our insights to propose BARE, a practical synthetic data generation method combining the diversity of base models with the quality of instruct models.
-BARE uses a base model to generate an initial set of diverse but potentially lower quality data, after which an instruct-tuned model individually refines each example from the initial set – as shown in the example in Figure 6.
-In the base step, the base model uses a prompt with minimal few-shot examples to generate.
-In the refine step, the instruct-tuned model is instructed to retain the concept of the base model generation while improving it according to specific criteria (e.g., realism, correctness).
-This retains the overall diversity of the base model generated set while exerting greater control over the quality of the final generation.
-Importantly, few-shot examples are only required in the base step to ensure formatting (though they can also be included in the refine step), and only in small amounts.
-Thus, BARE is especially useful in domains with very limited real-world data as little human effort is required to generate a diverse high-quality synthetic dataset.
-In our experiments, we limit ourselves to just three few-shot examples.
+Building on the relative advantages of base and instruction tuned models we introduce BARE (Figure 5).
+BARE is a practical few-shot synthetic data generation method that combines the diversity of base models with the fidelity of instruction tuned models.
+BARE uses a base model to generate an initial set of diverse but potentially lower quality data from very little seed data.
+An instruction-tuned model then individually refines each example from the initial set, improving it according to specific criteria (e.g., realism, correctness) while keeping the original concept.
+The refinement retains the overall diversity of the initial set while exerting greater control over the quality of the final generations.
+Importantly, diversity is elicited from the inherent properties of the base model rather than a large initial seed set, allowing for greatly improved data efficiency.
+Seed examples are only necessary to ensure the base model follows formatting (though they can also be included in the refine step).
+In our experiments, we use just three few-shot examples.
 In addition, we intentionally use very general prompts for BARE to demonstrate its flexibility, underscoring the potential for even greater improvement with tailored prompts.
-Representative prompts are included in Appendix B.
+Representative prompts are included in Appendix D.
 5 Evaluation
-We evaluate BARE for diversity, data quality, and downstream utility on the same domains and against the same baseline methods presented in Section 3.
-We implement BARE with Llama-3.1-70B-Base in the base generation stage and Llama-3.1-70B-Instruct in the refinement stage, and perform additional experiments with the Llama 3.1 8B family and with Llama 3.1 8/70B as the base model and GPT-4o as the refiner (Dubey et al., 2024; OpenAI, 2024c).
-We follow generally the same sampling temperatures as in Section 3.
-However, for Enron and Newsgroups, we instead sample from Llama-3.1-8B-Instruct at temperature $0.7$ in order to maintain coherence in generations.
-We always sample the refine model at a temperature of $0.7$.
-For evaluation of downstream utility for the generative tasks, we LoRA (Hu et al., 2021) fine-tune Llama-3.1-8B-Instruct for 4 epochs using the generated data, except for GSM8K where Llama-3.2-1B-Instruct (Dubey et al., 2024) is fine-tuned instead due to high baseline performance of the 8B model.
-Other fine-tuning hyperparameters are in Section A.1.1.
+We evaluate BARE on diversity, data quality, and downstream utility across the same domains and baselines presented in Section 3.
+BARE uses Llama-3.1-70B-Base for generation and Llama-3.1-70B-Instruct for refinement.
+We also experiment with the Llama 3.1 8B family and GPT-4o as the refiner (Dubey et al., 2024; OpenAI, 2024c).
+Sampling details are in Appendix A.
+To evaluate the downstream utility of BARE, we LoRA (Hu et al., 2021) fine-tune Llama-3.1-8B-Instruct for 4 epochs using the generated data, except for GSM8K where Llama-3.2-1B-Instruct (Dubey et al., 2024) is fine-tuned instead due to high baseline performance of the 8B model.
+Other fine-tuning hyperparameters are in Section B.1.1.
 The fine-tuned models are evaluated on a static $n=100$ test set for HotpotQA and PubMedQA, a static $n=500$ test set for GSM8K, and the full $n=442$ test set for LCB TOP.
-In this section, we will focus only on the generative tasks; the downstream evaluation process for classification tasks is presented in Section A.1.2 and detailed results for all domains can be found in Section A.2.
-Figure 7: Distribution of pairwise embedding similarities for Llama-3.1-70B-Base and Llama 3.1 70B BARE generations for GSM8K (top) and LCB TOP (bottom).
-The Base and BARE distributions are extremely similar for both tasks, indicating that refinement retains the diversity of the base generations.
-Note that frequency is of the cosine similarity score.
+In this section, we will focus only on the generative tasks; the downstream evaluation process for classification tasks is presented in Section B.1.2 and detailed results for all domains can be found in Section B.2.
+Figure 6: Distribution of pairwise embedding cosine similarity scores for Llama-3.1-70B-Base and Llama 3.1 70B BARE generations for GSM8K (top) and LCB TOP (bottom).
+The distributions are extremely similar for both tasks, indicating that refinement retains the diversity of base generations.
+Figure 7: Indistinguishability Rate for data generated by Llama-3.1-70B-Base and BARE across various datasets. BARE consistently improves the quality of data generated from base models.
 BARE Quality & Diversity.
-We begin by looking at the quality/diversity trends of BARE when compared with previous methods.
-From the histograms plotted in Figure 7, we can see that BARE effectively does not change the similarity distribution of generated data when compared to the base model at all – it is able to successfully retain the diversity of base generations.
-Detailed results with average embedding similarity scores can be found in Section A.2.
-Figure 8: Indistinguishability Rate for Llama-3.1-70B-Base and BARE data generation across various datasets. BARE consistently improves the quality of data generated from base models.
-Simultaneously though, looking at Figure 8, we can see that while retaining diversity, BARE leads to a monotonic increase in the IR for every domain - suggesting that it is able to lift the quality of the generations to be on par or in some cases even surpass directly sampling from an instruct model.
-Combined, this indicates that BARE is capable of leveraging the diversity of base models and quality of instruct-tuned models in its end generations.
-Figure 9: Fine-tuned model accuracy results on HotpotQA, PubMedQA, GSM8K, and LCB TOP for prompting methods using GPT-4o.
-Prompting methods have mixed effects on downstream performance when compared against standard sampling, while using GPT-4o with simply the Llama-3.1-8B-Base model in BARE improves consistently over all prompting methods.
-Figure 10: Accuracy of a Llama-3.1-8B-Instruct Model (HotpotQA, PubMedQA, LCB TOP) and Llama-3.2-1B-Instruct Model (GSM8K) finetuned on synthetic data generated using Base, Instruct, and BARE methods.
-Note that GSM8K 8B Instruct results are not shown as data generations were derailed.
+We begin by comparing the quality/diversity trends of BARE with other methods.
+From the histograms in Figure 7, we can see that BARE effectively does not change the similarity distribution of generated data when compared to the base model at all.
+Detailed results with average embedding similarity scores can be found in Section B.2.
+At the same time, we see from Figure 7 that BARE leads to a monotonic increase in the IR for every domain - suggesting that it is able to lift the quality of the generations to be on par or in some cases even surpass directly sampling from an instruct model.
+Combined, the IR and diversity measures indicate that BARE is capable of leveraging the diversity of base models and quality of instruction-tuned models in its end generations.
 Fine-tuned Model Accuracy.
-We now show the utility of BARE datasets as a whole.
-In Figure 10, we demonstrate the accuracy of a model fine-tuned on the datasets generated using different methods.
-Almost uniformly, BARE leads to greater improvements in downstream model quality than generating from either Base or Instruct models, and across all domains BARE-based data leads to the highest fine-tuned model accuracy.
-Impressively, in Figure 9, beyond just base and instruct sampling, we show that enhancing GPT-4o with BARE using only a small base model like Llama-3.1-8B-Base produces monotonically better fine-tuning data than all prompting methods discussed in Section 3.3 that use only GPT-4o.
-These clear results showcase BARE’s ability to out-perform the existing methods for high-quality, diverse data generation.
-Focusing on the RAFT domains, BARE improves upon the standard SOTA pipeline for fine-tuning LLMs for RAG by up to 18.4%, as seen with the 8B family on HotpotQA (standard RAFT is implemented in our Instruct generation results).
+We now show the utility of BARE-generated datasets as a whole.
+In Figure 8, we demonstrate the accuracy of a model fine-tuned on datasets generated using different methods.
+BARE-generated data leads to better downstream models than data from either Base or Instruct models almost uniformly, and across all domains training with BARE-generated data leads to the highest model accuracy.
+Surprisingly, in Figure 9, we find that BARE using the small Llama-3.1-8B base model to generate data and GPT-4o as a refiner, outperforms all advanced prompting methods discussed in Section 3.3 applied to GPT-4o directly.
+These clear results showcase BARE’s ability to out-perform existing methods for high-quality, diverse data generation in few-shot settings.
+Figure 8: Downstream accuracy of BARE compared to Llama-3.1 Base and Instruct models.
+Accuracy is measured on Llama-3.1-8B-Instruct Model (HotpotQA, PubMedQA, LCB TOP) and Llama-3.2-1B-Instruct Model (GSM8K) finetuned on synthetic data generated using Base, Instruct, and BARE methods.
+Note that GSM8K 8B Instruct results are not shown as generations derailed.
+In general, we find that BARE out-performs directly using the base and instruction fine-tuned models.
+Figure 9: Downstream accuracy for BARE versus other diversification methods.
+Fine-tuned model accuracy on HotpotQA, PubMedQA, GSM8K, and LCB TOP for data generated from prompting methods using GPT-4o.
+Prompting methods have mixed effects on downstream performance when compared against standard sampling.
+Alternatively, BARE using just with the Llama-3.1-8B Base as the generator and GPT-4o as the refiner significantly outperforms all over methods using GPT-4o.
+On RAFT domains, BARE improves upon the standard state-of-the-art pipeline for fine-tuning LLMs for RAG (RAFT) by up to 18.4%, as seen with the 8B family on HotpotQA (standard RAFT is represented in our Instruct generation results).
 BARE with both model families also outperforms existing RAFT pipelines on PubMedQA.
-While BARE with the Llama 3.1 70B family does not improve upon RAFT for HotpotQA, switching out the Llama-3.1-70B-Instruct refiner for GPT-4o does lead to an improvement (Section A.2).
 On GSM8K, BARE is the only method that provides useful training data.
-The un-trained model performance was 21.8%, and fine-tuning on BARE generated data achieves accuracies of 24.9% and 32.8% with the Llama 8B and 70B families, respectively.
+The un-trained model performance was 21.8%, and fine-tuning on BARE-generated data achieves accuracies of 24.9% and 32.8% with the Llama 8B and 70B families, respectively.
 Accuracy when training with data generated by single model methods either decreased accuracy or had little difference.
-In fact, training on data generated by Llama-3.1-70B-Instruct led to an accuracy of just 17.8%, which BARE with Llama-3.1-70B-Base refined by GPT-4o outperforms by 101% (35.8% accuracy).
+In fact, training on data generated by Llama-3.1-70B-Instruct led to an accuracy of just 17.8%, which BARE with Llama-3.1-70B-Base refined by GPT-4o outperforms by over 2×2$\times$2 × (35.8% accuracy).
 […]
-Similar to HotpotQA, on LCB TOP, BARE with the Llama 3.1 70B family shows less of an improvement than the 70B-instruct-only data generation method.
-Again, using GPT-4o as a refiner instead of Llama-3.1-70B-Instruct gives stronger results.
-This suggests that with the right refiner choice, BARE consistently provides gains.
-We show detailed results of refining with GPT-4o in all domains in Section A.2, emphasizing BARE’s consistently strong data generation capabilities.
-We also find that temperature ablations do not meaningfully affect the utility of instruct-tuned model generations, especially when compared to gains using BARE. For details, see Section A.3.
+We note that with the 70B family on both HotpotQA and LCB, BARE-generated data leads to less of an improvement than the instruct-generated data.
+In these cases, using GPT-4o as a refiner instead of Llama-3.1-70B-Instruct does lead to stronger results, indicating that the choice of refiner can strongly influence data quality: the Llama model was unable to sufficiently refine the base model generations, but the stronger GPT-4o was able to do so.
+We show detailed results of refining with GPT-4o in all domains in Section B.2, emphasizing BARE’s consistently strong data generation capabilities.
+For more detailed discussion, see Appendix C.
+Comparison to Real-World Data.
+To further show the quality of our synthetic data, we took the real GSM8K training data (the only domain of these 4 which has a real train set) and randomly sampled 1,000 examples from which to fine-tune the Llama-3.2-1B-Instruct model.
+The resulting trained model had a performance of $26.6\%$, surpassing the 8B/70B Base and Instruct models individually.
+Interestingly, though, it was surpassed by almost all BARE methods, most prominently when GPT-4o was used as a refiner with Llama-3.1-70B-Base (35.8%).
+While one might expect the real training data should be better than a synthetically generated dataset based on a small sample, the quality of synthetically generated data with an emphasis on quality and diversity can lead to better generalization then the original human generated training data.
+This is in line with other works’ findings for synthetic sets generated from large samples (Liu et al., 2023).
+Temperature & Scale Ablations.
+We also find that temperature ablations do not meaningfully affect the utility of instruction-tuned model generations, especially when compared to gains using BARE; for details, see Section B.3.
+Additionally, we found that BARE’s strong performance holds when scaling the amount of generated data; for details, see Section B.5.
+Figure 10: Accuracy at varying seed (few-shot) sizes of a Llama-3.2-1B-Instruct Model finetuned on synthetic data generated using Instruct and BARE methods on GSM8K.
+BARE performance is consistent throughout while Instruct improves with seed set size before converging with BARE.
 Instruct-Instruct Ablation.
-A possible challenge is that the gains in performance are due to a multi-step pipeline rather than combining a base and instruct-tuned model.
-To demonstrate the importance of using a base model in BARE, we performed an ablation where an instruct-tuned model’s generations are refined by an instruct-tuned model on the GSM8K task.
-The accuracy on the test set after switching Llama-3.1-70B-Base to Instruct in the first step drops from $29.8\%$ to $25.4\%$.
-This trend held when GPT-4o was the refiner as well, with a drop from $35.8\%$ to $30.8\%$.
-At the same time, the pattern of little change in the average pairwise similarity before and after refinement remains, indicating that diversity must be introduced in the first stage.
-Having earlier established that base models are most effective at diversity, we conclude that the use of base models is a necessary component of BARE.
-Detailed results from this ablation are in Section A.4.
+To verify BARE’s performance comes from the use of base models rather than the multi-step setup, we replace the base model in the first step of BARE with an instruction-tuned model on the GSM8K task.
+Switching Llama-3.1-70B-Base to Instruct drops the accuracy on the test set from $29.8\%$ to $25.4\%$ when refining with Llama-3.1-70B-Instruct and from $35.8\%$ to $30.8\%$ with GPT-4o.
+At the same time, the average pairwise similarity before and after refinement remains similar, indicating that diversity must be introduced in the first stage.
+Having established that base models are most effective at seed efficient diversity, we conclude that the use of base models is a key reason for the utility of BARE.
+Detailed results from this ablation are in Section B.4.
+Seed Data Scale Ablation.
+We additionally investigate the advantage of BARE compared to instruct-only methods as more seed data is made available.
+We evaluate results with BARE and Instruct using the Llama 3.1 70B family on GSM8K, ranging the seed set size from $3$ to $3000$ and continuing to generate 1,000 examples (Figure 10).
+Each generation randomly selects three examples from the seed set for the prompt.
+BARE consistently outperforms Instruct generation for small seed sets, highlighting the potential of BARE in making high-quality synthetic data generation easier to adopt and realistic in low-data domains.
+The methods converge and plateau for large seed set sizes, indicating that diversity sourced from the base model’s priors in BARE is always at least as effective as diversity from curated seed sets.
+In our experiments, convergence between BARE and Instruct generation occurred at around $100$ examples, or $10\%$ of the generated dataset size.
+We leave as future work the validation of the convergence point at larger dataset scales.
 6 Conclusion
-6.1 Key Contributions
-In summary, in this work we quantitatively investigated the quality and diversity of various synthetic data generation methods.
-Through this investigation, we find that base models are generally more diverse than instruct-tuned models while instruct-tuned models produce higher quality output than base models, validating hypotheses in the research community.
-These insights motivate the design of a better system to generate synthetic data, BARE.
+In this work we investigated the potential of using base models for few-shot synthetic data generation.
+Through this investigation, we find that base models are an effective source of diversity even when given minimal seed data, making them an attractive choice for data efficiency.
+These insights motivate the design of a system novelly leveraging base models in synthetic data generation, BARE.
 Through extensive experiments, we validate the importance of each step in BARE and demonstrate its ability to preserve base model diversity while enhancing output quality.
+The use and analysis of base models specifically in BARE opens an avenue for much future work as almost no prior work to our knowledge has studied their innate value (before post-training).
 Moreover, by fine-tuning on BARE-generated data for various domains, we underscore BARE’s practical utility, consistently outperforming existing synthetic data generation methods on downstream tasks such as GSM8K and LiveCodeBench, in addition to RAFT, for which we set a new SOTA.
-6.2 Future Work
-There are a lot of exciting directions for future work.
-For one, BARE is not necessarily the only way to elicit diversity in the synthetic data generation process.
-We were able to achieve all of our results with essentially basic prompting, leading us to believe there is room for even more improvement if one was to, for example, fine-tune the refiner specifically for this use case or use stronger refiner models.
-Introducing additional stages to BARE could also lead to improvements (e.g., additional refinement steps).
-Beyond multi-stage systems, the design space for diversity is vast – in one direction, training instruct-tuned models with different objective functions that encourage entropy is another area of exploration.
-Furthermore, the implications of a lack of high-quality, diverse generations from LLMs go beyond just synthetic data.
-For example, much of the work in inference time compute relies on models being able to generate sufficiently different possible trajectories and using these diverse trajectories (exploration) with a feedback signal to improve reasoning capabilities (Zhang et al., 2024a; Zelikman et al., 2022), yet only a few methods look explicitly at generator diversity (Wang et al., 2024).
-Lastly, the same BARE method demonstrated in this paper for synthetic training data can also be used with no modifications to generate synthetic evaluation sets, which are especially valuable in so many real-world domains with low-data availability.
-Acknowledgements
+Acknowledgments and Disclosure of Funding
 We would like to thank Justin Wong, Peter West, Raj Ammanabrolu, Alex Dimakis, Alex Xu, and Andrew Qin for their many insightful discussions about this work.
 Sky Computing Lab is supported by gifts from Accenture, AMD, Anyscale, Cisco, Google, IBM, Intel, Intesa Sanpaolo, Lambda, Microsoft, NVIDIA, Samsung SDS, SAP, and VMware.
+We would additionally like to acknowledge Databricks for their generous compute support.
 This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No.
@@
 Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
-Appendix A Additional Results
-A.1 Downstream Evaluation - Additional Details
-A.1.1 Fine-Tuning Task Hyperparameters
+Appendix A Sampling Details
+During data generation, all base models were sampled at a default temperature of $0.7$.
+For instruction-tuned-only generations we sampled at the highest temperature at which we experimentally found data generation to still be coherent, which is $1.0$ for Llama models and $1.2$ for GPT-4o.
+However, for Enron, we sample from GPT-4o at a temperature of $1.0$ to maintain generation coherence.
+Similarly, for Enron and Newsgroups, we sample for Llama-3.1-8B-Instruct at a temperature of $0.7$.
+Instruction-tuned models used during refinement are always sampled at a temperauter of $0.7$.
+Appendix B Additional Results
+B.1 Downstream Evaluation - Additional Details
+B.1.1 Fine-Tuning Task Hyperparameters
 We list below the fine-tuning hyperparameters that were used in common for HotpotQA, PubMedQA, GSM8K, and LCB TOP.
@@
   • LoRA Dropout: 0.0
-A.1.2 Classification Task Setup
-The generated data is used to train a BERT-based classifier [Devlin et al., 2018] for 2 epochs on Enron and 9 epochs on Newsgroups.
+B.1.2 Classification Task Setup
+The generated data is used to train a BERT-based classifier Devlin et al.
+[2018] for 2 epochs on Enron and 9 epochs on Newsgroups.
 The trained models are evaluated on a static test set with $n=500$ examples for each domain.
-A.2 Core Experiment Results - All Domains
+B.1.3 Compute
+We use 1-4 A100s for serving open-source models for data generation tasks (8B and 70B models) and up to a single node of 8xH100s for training purposes.
+LoRA fine-tuning takes on the order of minutes and data generation for 10,000 examples for an 70B model using 90 cores takes approximately an hour.
+B.2 Core Experiment Results - All Domains
 This appendix contains diversity, IR, and downstream performance results for all core experiments: generation with Llama 3.1 8B and 70B Base and Instruct models, BARE with Llama 3.1 models of both families, and BARE with the use of GPT-4o.
@@
 BARE GPT-4o + Llama 3.1 70B Base 0.356 88.5% 0.846
-Table 4: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on on a randomly selected static $n=500$ subset of Newsgroups.
+Table 4: Average pairwise embedding cosine similarity, IR, and downstream accuracy results on a randomly selected static $n=500$ subset of Newsgroups.
 A BERT model with a classification head was trained for $9$ epochs on the generated data $(n=500)$.
@@
 BARE GPT-4o + Llama 3.1 70B Base 0.471 68% 27.4%
-A.3 Independent Sampling Temperature Ablations - HotpotQA, PubMedQA, and LCB TOP
+B.3 Independent Sampling Temperature Ablations - HotpotQA, PubMedQA, and LCB TOP
 This appendix contains diversity, IR, and downstream performance results for our temperature ablation experiments.
@@
 BARE Llama 3.1 8B (All $t=0.7$) 0.462 47% 28.1%
-A.4 BARE First Stage Ablations - GSM8K
-This appendix contains diversity, IR, and downstream performance results for our ablation replacing the first stage of BARE with an instruct-tuned model, specifically Llama-3.1-70B-Instruct.
+B.4 BARE First Stage Ablations - GSM8K
+This appendix contains diversity, IR, and downstream performance results for our ablation replacing the first stage of BARE with an instruction-tuned model, specifically Llama-3.1-70B-Instruct.
 We refine using Llama-3.1-70B-Instruct and GPT-4o, and investigate the change in downstream performance compared to standard BARE (using Llama-3.1-70B-Base in the first stage).
@@
 BARE GPT-4o + Llama 3.1 70B Base 0.302 64% 35.8%
-Appendix B GSM8K Prompt Examples
+B.5 BARE Training Set Scale Ablations
+This appendix contains downstream performance results for our ablation where we scale up the number of BARE generated data points from $1,000$ to $10,000$ in an effort to investigate whether the trends hold at larger scales.
+We focused our investigation of this particular ablation on the Llama-3.1-8B model family, and across both the LCB TOP and PubMedQA domains we find that the BARE trends consistently hold at scale.
+Table 13: Downstream accuracy results on two evaluation sets (LCB and PubMedQA) using 10,000 BARE-generated data points and the Llama-3.1-8B model family.
+Models were fine-tuned for 2 epochs rather than the 4 used for 1,000 points due to compute considerations.
+Evaluation Set Base Instruct BARE
+LCB TOP 0.192 0.181 0.233
+PubMedQA 0.512 0.645 0.658
+Appendix C Additional Discussion
+In this appendix we further discuss our evaluation results and propose hypotheses for certain trends or observations.
+Why is BARE generation sometimes worse than instruction-tuned generation?
+As mentioned in the main paper, the choice of refiner plays an important role in the BARE pipeline.
+A model that is proficient at generating data may not be proficient at improving data.
+Mitigation of this issue can include tailoring prompts for specific models, but we chose to keep our prompts consistent and general to focus on the core idea of leveraging base mode generations.
+Why is instruction-tuned generation sometimes worse than base generation?
+A key insight to our work is the diversity-quality tradeoff between base models and instruction-tuned models. In some cases, the additional diversity from base models outweighs the lower quality.
+Why do smaller models sometimes produce better data than larger models?
+The diversity-quality tradeoff between smaller and larger models is not well understood.
+While it is generally accepted that larger models are of higher quality, it is unclear whether they are similar in terms of diversity.
+In some cases, the smaller model exhibits higher diversity in its generations, allowing the resulting datasets to be of more utility despite the lower quality.
+Appendix D GSM8K Prompt Examples
 In this appendix, we provide exact prompts used for the GSM8K domains, representative of those used throughout this work.
@@
 Start each example with the question.
-Delimit the end of an example with the phrase ”END OF EXAMPLE” (all caps) on a new line.
+Delimit the end of an example with the phrase "END OF EXAMPLE" (all caps) on a new line.
 Persona Prompt {persona_description} Provide an example of a grade school math word problem that requires performing a sequence of elementary calculations using basic arithmetic operations.